OpenStudy (anonymous):

simplify sinx+cosx=(sinx/1-(cosx/sinx))+(cosx/1-(sinx/cosx))

myininaya (myininaya):

You mean prove the identity?

OpenStudy (anonymous):

Are you sure about the parts which say sinx/1 and cosx/1? Aren't they just sinx and cosx respectively?

OpenStudy (anonymous):

verify that it is an identity

myininaya (myininaya):

\[\sin(x)+\cos(x)=\frac{\sin(x)}{1-\frac{\cos(x)}{\sin(x)}}+\frac{\cos(x)}{1-\frac{\sin(x)}{\cos(x)}}\]?

OpenStudy (anonymous):

yes

myininaya (myininaya):

I would take the ugly side and tried to show it is the prettier side

myininaya (myininaya):

Get rid of the compound fractions on the right hand side

OpenStudy (anonymous):

I've tried that and i can't get them to equal out

myininaya (myininaya):

You have two terms. Both terms are fractions. Both fractions contain fractions. (compound fractions they are) Multiply both fractions be a certain 1 that will clear the "compoundedness".

OpenStudy (anonymous):

what?

myininaya (myininaya):

Do you have a question? I'm asking you to make the compound fractions into regular fractions.

OpenStudy (anonymous):

i did that

myininaya (myininaya):

\[\frac{\sin(x)}{1-\frac{\cos(x)}{\sin(x)}} \cdot 1+\frac{\cos(x)}{1-\frac{\sin(x)}{\cos(x)}} \cdot 1 \] What kinda 1 should you choose for the first 1? What kinda 1 should you choose for the second 1?

myininaya (myininaya):

Can you answer these questions?

OpenStudy (anonymous):

ya

myininaya (myininaya):

Ok so what did you choose for the 1st 1 and 2nd 1?

OpenStudy (anonymous):

\[\frac{ \sin^2x }{1-cosx ? }\]

myininaya (myininaya):

Ok I will tell you this step For the 1st 1 I chose the 1 to be sin(x)/sin(x) For the 2nd 1 I chose the 1 to be cos(x)/cos(x) So like this we have: \[\frac{\sin(x)}{1-\frac{\cos(x)}{\sin(x)}} \cdot \frac{\sin(x)}{\sin(x)}+\frac{\cos(x)}{1-\frac{\sin(x)}{\cos(x)}} \cdot \frac{\cos(x)}{\cos(x)}\]

myininaya (myininaya):

So we have: \[\frac{\sin(x) \cdot \sin(x)}{(1-\frac{\cos(x)}{\sin(x)})(\sin(x))}+\frac{\cos(x) \cdot \cos(x)}{(1-\frac{\sin(x)}{\cos(x)} ) (\cos(x))}\]

myininaya (myininaya):

Recall the distributive property a(b+c)=ab+ac Use this here on the bottoms for sure And you should be able to multiply the tops with no problems

myininaya (myininaya):

@karr Where did you go? :(

OpenStudy (anonymous):

sorry. okay i got that.. the top of the fractions equal 1 now....

OpenStudy (anonymous):

??